OpenStudy (erniellyd28):

given f(x)equals 3x minus 4 find f(4)

OpenStudy (anonymous):

f(x) is a fancy way of saying "y" f(4) is a fancy way of saying the "y"' value when x=4 f(x)=y=3x-4 f(4)=y=3(4)-4=8

OpenStudy (erniellyd28):

thank you so very much i dont know why but for its complicated